# PreLoRA: Hybrid Pre-training of Vision Transformers with Full Training and Low-Rank Adapters

### v1

@@ 1 Introduction
In recent years, the field of machine learning has undergone significant advances in the development and deployment of large models with ever-increasing capacity across diverse domains [14, 21, 19]. The notion of large models has evolved from architectures with millions to those comprising billions of parameters, enabling significant improvement in performance and generalization [13]. However, the massive parameter count that strengthens model capacity in turn introduces substantial computational and memory bottlenecks, rendering training prohibitively expensive [22, 25].
Several parameter-efficient approaches have been explored to reduce the number of trainable parameters while preserving model performance and generalization. Adapter methods [10] insert small trainable modules into each layer and keep the original parameters frozen. While this approach does reduce the number of trainable parameters, it introduces additional inference latency. The prefix [16] and the prompt tuning [15] techniques enable lightweight adaptation by optimizing a small set of learnable task-specific vectors that are added to the input sequence or intermediate activations. However, these methods often demonstrate limited expressivity when applied to complex tasks [24]. Recently, Low-Rank Adaptation (LoRA) [11] has gained widespread popularity because it leverages the flexibility of adapter methods with better expressivity than prompt-based methods, while maintaining minimal resource overhead. This is achieved by injecting low-rank trainable matrices into existing weight updates without adding any additional layers. While prior applications of LoRA have predominantly focused on fine-tuning, its integration into pretraining has the potential to accelerate training, reduce resource demands, and maintain model capacity.
@@
Motivation:
During the initial training epochs, the model undergoes substantial weight updates; however, as training progresses, the intrinsic rank of the updates decreases[11]. Figure 1 illustrates this behavior for a representative large Vision Transformer model. Plot (a) on the left shows that the weight norm variations are pronounced across all the modules during the initial epochs but gradually stabilize in the later half of the training loop. Despite the stabilization of weight norms, the model loss (shown in plot (b)) continues to reduce significantly. Therefore, transitioning to LoRA at this stage has the potential to effectively capture the remaining low-rank variations in weight norms.
Challenges:
Identifying the optimal transition point from full-parameter to LoRA-based training is non-trivial. An early transition can yield substantial performance gains, but often at the expense of accuracy. In contrast, a late transition preserves accuracy, but limits potential reductions in training time. The central challenge, therefore, lies in identifying a dynamic switching strategy that balances model accuracy and training performance. Enabling explicit user control over this switching point allows them to choose an accuracy-speed trade-off, making the model adaptable to specific use-case requirements. Moreover, it is evident from Figure 1 that although weight norms across modules exhibit broadly consistent trends, their magnitudes differ considerably. Consequently, applying a uniform low rank across all modules may inadequately capture module-specific convergence dynamics.
Contributions:
In this work, we present PreLoRA, a framework for integrating Low-Rank Adaptation (LoRA) into the pretraining of large-scale models and address the challenges mentioned above. We use Vision Transformer (ViT-Large) [5] with 300M parameters as a representative model for our experimental study. Our main contributions are as follows:
@@
    We propose a method for identifying the optimal switching point from full-parameter to LoRA-based training via a custom convergence test (Section 3.1) that incorporates loss and weight-change thresholds, enabling explicit control over the trade-off between training efficiency and model accuracy.
@@
    We develop a dynamic rank-assignment algorithm (Section 3.2) that assigns module-specific ranks based on its weight convergence at the point of transition, facilitating efficient utilization of model capacity.
@@
    We perform extensive experiments on ViT-Large (Section 4.2), to validate the proposed approach, with a focus on how key design choices like the convergence criteria and warmup window size impact training efficiency and final model performance.
  4. 4.
    We present a generalizable framework for applying parameter-efficient techniques during pretraining, enabling faster and more resource-efficient training of large models across diverse domains.
@@ 2 Related Work
In an earlier work, the adapter method [9] popularized the domain of parameter-efficient fine-tuning by introducing lightweight, trainable modules inserted into pretrained models, thereby minimizing the number of parameters updated during adaptation. Since its inception, various adapter variants have been developed and applied across diverse domains, including NLP, CV and speech recognition [9, 1, 8, 20]. In parallel, prompt-tuning [15] and prefix-tuning [16] emerged as alternative approaches, which reduce trainable parameters by appending learnable prompts or prefixes to the input sequence, rather than modifying the model’s architecture. These methods have gained widespread adoption for fine-tuning LLMs [26, 2, 27] and vision models [30, 29, 12], offering flexibility in scenarios with limited computational resources. Our work, however, centers on Low-Rank Adaptation (LoRA) [11], which provides a third paradigm by decomposing weight updates into low-rank matrices, thus achieving parameter efficiency without introducing new modules. LoRA stands out for its superior efficiency, stronger expressivity, and ease of integration, as evidenced by empirical studies [24, 4] compared to adapters and prompt-tuning. Its applications span language models, vision transformers, multimodal systems, and even reinforcement learning frameworks [7, 17, 28, 18, 6, 4], with extensions like QLoRA further combining quantization for enhanced memory savings during fine-tuning.
Despite these advancements, most PEFT methods, including LoRA, have been predominantly explored in the context of fine-tuning pretrained models rather than full pretraining from scratch. This focus creates a notable gap, as pretraining large models demands immense computational resources, and efficient methods could democratize access to custom foundation models. Our work aims to bridge this gap by investigating LoRA’s efficacy in pretraining scenarios, potentially enabling scalable training on resource-constrained hardware. A recent study [3] applied LoRA to train a spatiotemporal attention-based model, accelerating it on processing-in-memory (PIM) architectures. Their approach monitors weight convergence during training and dynamically switches to LoRA upon detection, reducing overall training costs. However, it requires running two model copies in parallel (the full model and the LoRA variant) and performs a t-test on their losses for convergence checking. This dual-model setup is particularly inefficient for large models, as it doubles memory usage, hampers scalability, and introduces overhead that could be mitigated through more lightweight monitoring techniques, such as periodic loss sampling. Future directions might include hybrid PEFT strategies that combine LoRA with quantization or distillation to further optimize pretraining efficiency.
@@ 4 Experimental Results
Figure 4: Ablation Study: Comparison of different PreLoRA settings to the full baseline model in terms of accuracy (a,c,d) and speed (b). Only epoch 100 and later are shown to highlight the part of the training cycle that correspond to different hyperparameter settings.
@@ 4.1 Experimental Setup
Model and Data:
To showcase the advantages of our PreLoRA framework, we employ the ViT-Large architecture as a representative vision model, trained from scratch on the ImageNet-1k dataset. For an effective configuration to train ViT-Large from scratch on ImageNet-1K dataset, we adopt the training configuration prescribed by Steiner et al. [23].
Hardware:
Training was conducted in a distributed setting on a supercomputing cluster. A total of 64 NVIDIA A100 GPUs were used across 16 compute nodes connected via NVLink, to enable high-bandwidth and low-latency inter-GPU communication. Each GPU features 6912 CUDA cores and 432 Tensor cores, supporting FP32, FP16, BF16, and Int8 precisions. They have 40 GiB of HBM2, offering a memory bandwidth of 6.4 TB/s and thermal design power of 600 watts. The host is equipped with a 2.8 GHz AMD EPYC Milan 7543P 32-core CPU and 512 GB of DDR4 RAM.
Software and Code:
The HuggingFace library (v0.33.0) provided the model architecture, which was implemented in PyTorch (v2.7.1) for distributed training with LoRA support from PEFT (v0.15.2). The project is open-source, with code publicly available on GitHub: https://anonymous.4open.science/r/PreLoRA-7324, which also contains the data processing and visualization scripts using multiple Python packages.
Hyperparameters:
In our PreLoRA approach, following Dahal et al. [3] we employed 3 consecutive windows of 3 epochs each for the partial convergence test (Algorithm 1) by setting kk=3 and mm=3. Testing for convergence too frequently risks false positives from local minima, whereas overly large windows can prolong full-parameter training unnecessarily. We define $\alpha$ = [query (q), key (k), value (v), dense (d), output (o)] as the set of target modules for training in the PreLoRA framework. Our experiments evaluate how the hyperparameters $\tau$ and $\zeta$, which control the strictness of the convergence test, mediate the trade-off between accuracy and efficiency. We further investigate the effect of the warmup window size $w$, which dictates how many epochs the LoRA modules are trained jointly with the full base model prior to freezing its weights. For the dynamic rank assignment algorithm, ranks were constrained between 8 and 64.
@@ 4.2 Results
We demonstrate the effectiveness of PreLoRA by comparing its performance across a range of hyperparameter settings against the full model baseline. To this end, Section 4.2.1 evaluates two key design choices: the strictness of the partial convergence test, determined by $\tau$ and $\zeta$, and the warmup window size $w$, which dictates how long the LoRA modules train alongside the full model before freezing its weights. Section 4.2.2 follows, presenting results that provide a broad overview of the performance gains achieved with PreLoRA in terms of time, throughput, and memory.
@@ 4.2.1 Accuracy-Efficiency Trade-offs
Strictness of partial convergence test ($\tau,\zeta$): As discussed previously in Section 3.1, determining an appropriate transition point from full-parameter training to LoRA-based training is a crucial component of the PreLoRA framework. In our method, this transition is guided by a convergence criterion defined through changes in weight norms ($\tau$) and training losses ($\zeta$). Together, these indicators capture partial convergence and trigger the switch. Table 1 summarizes different configurations of these constraints, ranging from a relaxed setting (Exp1) to a strict setting (Exp3).
Experiment $\tau(\%)$ $\zeta(\%)$
Exp1 1.00 5.00
Exp2 0.50 2.50
Exp3 0.25 1.00
Table 1: PreLoRA: experimental settings for $(\tau,\zeta)$ thresholds.
@@
A relaxed constraint causes the partial convergence criterion to be satisfied earlier, which is expected to yield substantial improvements in training efficiency, albeit at the cost of reduced accuracy, and vice versa. Figure 4(b) shows that the most relaxed setup (Exp1) accelerates training by approximately 40%, whereas the performance gain from the strictest setting (Exp3) is around 28%. This gain, however, comes with a trade-off: the training cross-entropy loss (shown in Figure 4(a)) converges to around 2.0 in Exp1, compared to 1.75 in Exp3. A similar trend is observed in training and validation accuracy (Figure 4). Interestingly, despite the marginally degraded accuracy in terms of other metrics, Exp1 attains a higher top-1 validation accuracy than the other experiments. Therefore, for tasks where accuracy is paramount, switching later is advisable, whereas for applications that emphasize faster training and can tolerate a modest accuracy drop, an earlier switch is preferable. A significant observation is that, even during the pretraining phase, LoRA adaptation can capture the learning trend, despite a substantial reduction in the learnable parameters (from 300M to 30M).
Effect of warmup window size ($w$): In this set of experiments, we vary the hyperparamter $w$ that controls the number of epochs (after passing the partial convergence test) we let the LoRA parameters train along with the full model before freezing the latter. The objective is to study the effect of the guidance provided by the full model to the LoRA parameters for different window sizes. The set of values we use are w=5,10,15w={5,10,15}. We set $\tau=0.50$ and $\zeta=2.5$ corresponding to Exp2 from Table 1. This setting results in the training mode switching to only the LoRA parameters (with the base model frozen) at around the $150$-th epoch. We know from Figure 1 that it is during this phase of the training cycle that the base model weights largely stabilize for it to be able to provide the necessary guidance. The results are shown in Figure 5.
@@
Figure 5: Ablation Study: (a) Cross-Entropy loss comparison and (b) epoch speedup, for different LoRA warmup windows.
@@
Figure 6: Effect of warmup window size on weight norms: Weight norms (Query) during warmup windows for (a) the Base model and (b) LoRA parameters.
@@
We present the loss curves (Figure 5(a)) and epoch times (Figure 5(b)) for warmup window sizes $w\in{5,10,15}$, with the full model serving as the baseline, evaluated from epoch 135 onward. The epoch times exhibit the expected trend: a shorter warmup window triggers an earlier transition to LoRA-only training, thereby yielding greater performance gains.
For cross-entropy loss, the baseline model achieves marginally better performance than the PreLoRA variants. Notably, the configuration with $w=15$ temporarily outperforms the baseline during the warmup phase (highlighted between the vertical lines in Figure 5(a)), before converging to the performance of the other window sizes. We hypothesize that excessively short warmup windows fail to provide sufficient guidance for adapting the LoRA parameters, while overly long warmup windows allow the base model to absorb the majority of weight updates, limiting adaptation in the LoRA parameters. Consequently, after the warmup phase, the $w=15$ configuration exhibits a momentary increase in loss until the LoRA weights stabilize, eventually converging to the performance of other settings.
To further investigate this hypothesis, we analyze the weight norm dynamics of the base and LoRA parameters. As shown in Figure 6(a), a larger warmup window ($w=15$) results in higher weight norms in the base model, whereas Figure 6(b) demonstrates the opposite effect for the LoRA parameters, with larger windows corresponding to smaller weight norms. In summary, longer warmup windows provide limited accuracy benefits while delaying performance improvements. Among the tested configurations, $w=10$ achieves the most favorable trade-off between training efficiency and final model performance.
4.2.2 Time, compute, and memory
@@ 5 Conclusion
In summary, this work proposes a framework that leverages LoRA to reduce the number of trainable parameters during pre-training of a large vision model. It provides a methodology to identify a transition point between full-parameter and low-parameter training, enabling explicit control over the accuracy–speed trade-off. While evaluated on ViT-Large, the approach is generalizable to other deep learning architectures. As future work, we plan to investigate online hyperparameter optimization techniques to establish a more principled and generalizable method for defining the convergence criterion in the partial convergence test, thereby enhancing the robustness, scalability, and applicability of the approach across diverse domains.
Acknowledgment

### v2

# Argonne National Laboratory2, Lemont, IL, USA Red Hat3, Raleigh, NC, USA

@@ 1 Introduction
Recent advances in machine learning have led to the development and deployment of large models across diverse domains [14, 22, 20]. Their ever increasing size from millions to billions of parameters while enabling significant improvement in performance and generalization [13], has introduced substantial computational and memory bottlenecks, rendering training prohibitively expensive [23, 26].
Parameter-efficient approaches have been explored to reduce trainable parameters while preserving model performance and generalization. Adapter methods [10] insert small trainable modules into each layer with frozen base weights. However, this reduction comes at the cost of increased inference latency. Prefix and prompt tuning [16, 15] techniques optimize a small set of learnable task-specific vectors that are added to the input sequence or intermediate activations, but often exhibit limited expressivity when applied to complex tasks [25]. Low-Rank Adaptation (LoRA) [11] combines the flexibility of adapter methods with better expressivity than prompt-based methods, while maintaining minimal resource overhead by injecting low-rank trainable matrices into existing weight updates without adding any additional layers. While predominantly applied to fine-tuning, LoRA’s integration into pretraining has the potential to accelerate training, reduce resource demands, and maintain model capacity.
@@
Motivation: During the initial training epochs, the model undergoes substantial weight updates, which gradually stabilize as training progresses. For a representative large Vision Transformer model, Plot (a) of Figure 1 shows pronounced weight norm variations across all the modules initially followed by stabilization in the later epochs, while the model loss (shown in plot (b)) continues to decrease. These observations suggest that, in later epochs, weight updates primarily reflect small refinements rather than large exploratory changes across all parameters. We hypothesize that the low-rank constraint of LoRA is sufficient to capture these refinements without significant accuracy loss.
Challenges: Identifying the optimal point to transition to LoRA-based training is non-trivial: an early transition can speed training but reduce accuracy, whereas a late transition preserves accuracy, but limits performance gains. A dynamic switching strategy, enabling explicit user control over this switching point, balances this accuracy–speed trade-off. Additionally, Figure 1 shows that while weight norms broadly exhibit consistent trends across modules, their magnitudes might differ considerably, indicating that a uniform low rank assignment may fail to capture module-specific convergence dynamics.
Contributions: In this work, we explore the integration of Low-Rank Adaptation (LoRA) into the pretraining of large-scale models. To that end, we have developed a framework, which we named PreLoRA that aims to address the challenges mentioned above. We use Vision Transformer (ViT-Large) [5] with 300M parameters as a representative model for our experimental study. Our main contributions are as follows:
@@
    We propose a custom convergence test (Section 3.1) that identifies the switching point from full-parameter to LoRA-based training using loss and weight-change thresholds and provides users explicit control over the accuracy–speed trade-off.
@@
    We introduce a dynamic rank-assignment algorithm (Section 3.2) that assigns module-specific ranks at the transition point, improving parameter efficiency.
@@
    We conduct experiments on ViT-Large (Section 4.2), to study how key design choices like the convergence criteria and warmup window size impact training efficiency and final model performance.
These contributions form a generalizable framework for applying parameter-efficient techniques during pretraining, enabling faster and more resource-efficient training of large models across diverse domains. We further extend this framework by integrating an existing LoRA-based technique and present a targeted study demonstrating the complementary performance improvements it provides.
@@ 2 Related Work
Early work on adapters [9] established parameter-efficient fine-tuning (PEFT) by introducing lightweight, trainable modules into pretrained models, reducing the number of trainable parameters during adaptation. Numerous adapter variants have since been explored across NLP, computer vision, and speech recognition [9, 1, 8, 21]. In parallel, prompt-tuning [15] and prefix-tuning [16] emerged as architecture-preserving alternatives that append learnable prompts or prefixes to the input sequence, and have been widely adopted for fine-tuning large language and vision models under resource constraints [27, 2, 28, 31, 30, 12]. Our work instead centers on Low-Rank Adaptation (LoRA) [11], which provides a third paradigm by decomposing weight updates into low-rank matrices, thus achieving parameter efficiency without introducing new modules. LoRA has demonstrated superior efficiency, expressivity, and ease of integration compared to adapters and prompt-tuning [25, 4], across multiple application domains [7, 18, 29, 19, 6, 4].
Despite these advances, PEFT methods, including LoRA, have been studied primarily for fine-tuning, with only a small number of works having investigated the use of LoRA during pretraining from scratch. One work [3] applies LoRA to a spatiotemporal attention-based model on processing-in-memory (PIM) architectures by dynamically switching to a low-rank parameterization based on weight convergence. While this reduces training cost, it requires maintaining two model instances in parallel–a full model and its LoRA counterpart–and relies on statistical tests over their losses to detect convergence. This substantially increases memory usage and overhead, limiting scalability. Another work, ReLoRA [17] provides an alternative way of incorporating low-rank updates during pre-training. After an initial warmup phase, the model transitions to a lower-rank parameterization and periodically merges and reinitializes low-rank parameters to emulate full-rank optimization. However, ReLoRA applies a fixed and pre-determined uniform rank across all layers and modules throughout the whole pretraining run. This approach does not account for variation in convergence dynamics across layers and modules, which flexible, layer- or module-specific rank assignments could better capture, potentially enabling more efficient parameter allocation and improved performance.
@@ 4.1 Experimental Setup
Model and Data:To showcase the advantages of our PreLoRA framework, we employ the ViT-Large architecture as a representative vision model, trained from scratch on the ImageNet-1k dataset. For an effective configuration to train ViT-Large from scratch on ImageNet-1K dataset, we adopt the training configuration prescribed by Steiner et al. [24].
Hardware: Training was performed on a distributed supercomputing cluster using 64 NVIDIA A100 GPUs across 16 NVLink-connected nodes for high-bandwidth, low-latency communication. Each GPU provides 40 GB HBM2 memory with support for FP32, FP16, BF16, and INT8 precision. Compute nodes are equipped with 2.8 GHz AMD EPYC Milan 7543P (32-core) CPUs and 512 GB DDR4 RAM.
Software and Code: Experiments were implemented in PyTorch (v2.7.1) using HuggingFace Transformers (v0.33.0) with LoRA support via PEFT (v0.15.2). The open-source code, including data processing and visualization scripts, is available at https://anonymous.4open.science/r/PreLoRA-D624.
[…]
@@ 4.2 Results
(a) CE Loss
(b) Training Time
(c) Top-1 Train Accuracy
(d) Top-1 Validation Accuracy
Figure 4: Ablation Study. Comparison of different PreLoRA settings to the full baseline model in terms of accuracy (a,c,d) and time (b). Displays epoch 100 and later for clarity
We demonstrate the effectiveness of PreLoRA by comparing its performance across a range of hyperparameter settings against the full model baseline. To this end, this section evaluates two key design choices: the strictness of the partial convergence test, determined by $\tau$ and $\zeta$, and the warmup window size $w$, which dictates how long the LoRA modules train alongside the full model before freezing its weights. This is followed by results that provide a broad overview of the performance gains achieved with PreLoRA in terms of time, throughput, and memory. Finally, we integrate ReLoRA’s merge-and-reinitialize technique to achieve further performance gains.
@@ 4.2.1 Accuracy-Efficiency Trade-offs
Strictness of partial convergence test ($\tau,\zeta$): As discussed previously in Section 3.1, determining an appropriate transition point from full-parameter training to LoRA-based training is a crucial component of the PreLoRA framework. In our method, this transition is guided by a convergence criterion defined through changes in weight norms ($\tau$) and training losses ($\zeta$). Together, these indicators capture partial convergence and trigger the switch. We evaluate three partial convergence settings ranging from relaxed to strict thresholds: Exp1 $(\tau{=}1.0\%,,\zeta{=}5.0\%)$, Exp2 $(\tau{=}0.5\%,,\zeta{=}2.5\%)$, and Exp3 $(\tau{=}0.25\%,,\zeta{=}1.0\%)$.
@@
A relaxed constraint causes the partial convergence criterion to be satisfied earlier, which is expected to yield substantial improvements in training efficiency, albeit at the cost of reduced accuracy, and vice versa. Figure 4(b) shows that the most relaxed setup (Exp1) accelerates training by approximately 40%, whereas the performance gain from the strictest setting (Exp3) is around 28%. This gain, however, comes with a trade-off: the training cross-entropy loss (shown in Figure 4(a)) converges to around 2.0 in Exp1, compared to 1.75 in Exp3. A similar trend is observed in training and validation accuracy (Figure 4).Despite the marginally degraded accuracy in terms of other metrics, Exp1 attains a higher top-1 validation accuracy than the other experiments. Therefore, for tasks where accuracy is paramount, switching later is advisable, whereas for applications that emphasize faster training and can tolerate a modest accuracy drop, an earlier switch is preferable. A significant observation is that, even during the pretraining phase, LoRA adaptation can capture the learning trend, despite a substantial reduction in the learnable parameters from 300M to 30M.
(a) Cross-Entropy Loss
(b) Epoch Speedup
@@
Figure 5: Ablation Study. Comparison of different LoRA warmup windows: (a) Cross-Entropy loss, and (b) training speedup.
4.2.2 Effect of warmup window size ($w$):
In this set of experiments, we vary the hyperparameter $w$ that controls the number of epochs (after passing the partial convergence test) we let the LoRA parameters train along with the full model before freezing the latter. The objective is to study the effect of the guidance provided by the full model to the LoRA parameters for different window sizes. The set of values we use are w=5,10,15w={5,10,15}. We set $\tau=0.50$ and $\zeta=2.5$ corresponding to Exp2. This setting results in the training mode switching to only the LoRA parameters (with the base model frozen) at around the $150$-th epoch. We know from Figure 1 that it is during this phase of the training cycle that the base model weights largely stabilize for it to be able to provide the necessary guidance. The results are shown in Figure 5. We present the loss curves (Figure 5(a)) and epoch times (Figure 5(b)) for warmup window sizes $w\in{5,10,15}$, with the full model serving as the baseline, evaluated from epoch 135 onward. The epoch times exhibit the expected trend: a shorter warmup window triggers an earlier transition to LoRA-only training, yielding greater performance gains.
@@ 4.2.2 Effect of warmup window size ($w$):
Figure 6: Effect of Weight norms (Query) on warmup windows for (a) the Base model and (b) LoRA parameters.
For cross-entropy loss, the baseline slightly outperforms the PreLoRA variants overall. The w=15 configuration briefly surpasses the baseline during the warmup phase (Figure 5(a)) before converging with other settings. Short warmup windows may insufficiently guide LoRA adaptation, whereas longer warmup periods shift most updates to the base model, limiting LoRA learning. Hence, w=15 shows a temporary post-warmup loss increase before stabilizing and converging.
To examine this, we analyzed weight norm dynamics of the base and LoRA parameters. Figure 6(a) shows that a larger warmup window (w=15) increases base model weight norms, while Figure 6(b) shows that LoRA norms decrease with longer windows. Overall, longer warmups offer limited accuracy gains and slower improvements, with w=10 providing the best trade-off between efficiency and final performance.
4.2.3 Time, compute, and memory:
@@
4.2.4 Integration with ReLoRA:
Figure 8: ReLoRA Integration: Time Per Epoch
PreLoRA reduces parameter count after partial convergence, typically in the later training stages. To explore whether earlier epochs can also benefit, we integrate the ReLoRA strategy [17] from the start, beginning with several full-parameter epochs before transitioning to periodic merge-and-reinitialize updates. After the first transition, the rank is reduced from 1024 to 256 (300M to 114M parameters), continuing until partial convergence. At this point our rank assignment algorithm sets layer-specific rank between 8 and 64, further reducing the parameter count to less than 10% of the original with training continuing at these ranks for the remainder of the run. This hybrid approach preserves validation accuracy with minimal impact on cross-entropy loss, improves training speed by $\sim$ 122 seconds per epoch (42% performance gain), as shown in Figure 8. All experiments use the Exp1 hyperparameters with a 10-epoch warmup following convergence detection. Epoch times show a temporary rise from epochs 200 to 210, corresponding to the 10-epoch warmup after partial convergence.
@@ 5 Conclusion
In summary, we present and evaluate a LoRA-based approach for reducing trainable parameters during large vision model pre-training. Our results show that, under appropriate convergence criteria, low-rank constraints are sufficient to capture late-stage training refinements without significant accuracy loss. The method identifies an effective transition strategy between full-parameter and low-parameter training, giving users explicit control over the accuracy–speed trade-off. Although evaluated on ViT-Large, the approach generalizes to other architectures and is further enhanced through integration with ReLoRA. Future work will explore online hyperparameter optimization to establish a principled convergence criterion and improve robustness and scalability across domains.
6 Acknowledgment